### v1

# ChatWise: AI-Powered Engaging Conversations for Enhancing Senior Cognitive Wellbeing

Abstract
Cognitive health in older adults presents a growing challenge. While conversational interventions show feasibility in improving cognitive wellness, human caregiver resources remain overburdened. AI-based methods have shown promise in providing conversational support, yet existing work is limited to implicit strategy while lacking multi-turn support tailored to seniors. We improve prior art with an LLM-driven chatbot named ChatWise  for older adults. It follows dual-level conversation reasoning at the inference phase to provide engaging companionship. ChatWise thrives in long-turn conversations, in contrast to conventional LLMs that primarily excel in short-turn exchanges. Grounded experiments show that ChatWise significantly enhances simulated users’ cognitive and emotional status, including those with Mild Cognitive Impairment.
ChatWise: AI-Powered Engaging Conversations
for Enhancing Senior Cognitive Wellbeing
Zhengbang Yang, Zhuangdi Zhu George Mason University, Fairfax, VA, USA {zyang30,zzhu24}@gmu.edu
1 Introduction
[…]
In this work, we aim to draw on the advance of AI to provide older adults with engaging conversational support that improves their cognitive function and reduces social loneliness, serving as accessible alternatives to human companions. Previous clinical studies aimed at socially isolated older adults showed the existence of causal relationships between interviewer strategy and interviewee response  (Cao et al., 2021), revealing that conversation behavior can have a measurable influence on interlocutors, which inspired us to develop principled yet efficient methods that transform clinical insights into strategic dialogue design leveraging the advanced reasoning ability of LLMs.
In response, we propose an LLM-driven chatbot named ChatWise. Unlike traditional chatbot driven with implicit conversational goals, ChatWise employs dual-layer conversation generation, which first derives categorized macro-level information, including user emotion states to suggest meta-conversational strategies, which then guides the micro-level utterance generation to improve both user engagement and cognitive outcomes over multi-turn dialogue interactions. Figure 1 overviews its development design. Our work provides multifold contributions:
  • ChatWise is a tuning-free yet principled framework for developing AI chatbots with enhanced multi-turn, daily-themed conversational support. Our dual-level conversation generation design integrates clinical insights into LLM reasoning and can be readily applied to various LLMs.
  • ChatWise follows grounded development and evaluation using digital twins (Hong et al., 2024), which are simulated users constructed by tuning LLMs on de-identified, real-world dialogues between seniors and professional caregivers of a clinical trial (I-CONECT, 2024). Interactions with digital twins demonstrated that ChatWise significantly enhances users’ engagement and cognitive status with multi-turn interactions, especially for users with Mild Cognitive Impairment (MCI) (Figure 2).
  • Our comparative studies demonstrated that integrating macro-level strategies into conversation generation is the key contributor to enhancing user engagement. Its gain on cognitive metrics remains consistent across different backbone LLMs. ChatWise’s advantages become more pronounced in longer conversations, in contrast to conventional LLMs that primarily excel in short-turn question-answer exchanges.
  • Our dialogue analysis revealed key empirical insights, including the pattern of predominant interaction strategies across users. While transient user emotional fluctuations were less tractable, multi-turn conversations with ChatWise have shown to notably improve user emotions over time. This highlights the value of long-term, principled conversational support in enhancing seniors’ cognitive and emotional well-being. We hope these findings can inspire future research, including RL-based LLM optimization.
Figure 1: Overview of ChatWise Development.
Figure 2: ChatWise notably improves multiple user cognitive metrics through 10-turn dialogues (Sec 4.2).
2 Related Work
AI-powered Chatbots for Seniors: Social interaction plays a crucial role in the cognitive health of seniors  (Patil and Braun, 2024; Fiori and Jager, 2012). AI-powered chatbots have emerged as a promising alternative  (Rodríguez-Martínez et al., 2023; Owan et al., 2023). Recent efforts span commercial products (sen, 2024; ell, 2024) and research prototypes focused on emotional support and audio assistance  (Yang et al., 2024; Liu et al., 2021b; sen, 2024; Ryu et al., 2020; Hong et al., 2024). Unlike prior work, our approach prioritizes user conversational engagement, which has been empirically shown to enhance both cognitive and emotional status.
Conversational Strategies: Liu et al. (2021a) curated a dataset with annotated strategies, demonstrating the effectiveness of Helping Skills Theory Hill (2020) in providing emotional support. Yuan et al. (2023) examined the causal relationships between dialogue acts (DAs) and participants’ emotional states in a clinical trial (I-CONECT, 2024), emphasizing the impact of strategic interventions in tele-mediated dialogues. Seo et al. (2021) identified key strategies for improving child patient-provider communication through semi-structured interviews. Few works have systematically integrated these strategies into the reasoning flow of AI chatbots. Our approach fills this gap by embedding structured conversational strategies for automatically enhancing user engagement.
Multi-turn chatbot exploration: Recent advances in AI dialogue systems predominantly focused on short-turn interactions  (Owan et al., 2023; Dam et al., 2024). While some pioneering research has explored multi-turn optimization through Reinforcement Learning (RL) approaches for LLMs (Verma et al., 2022; Zhou et al., 2024; Abdulhai et al., 2023; Gao et al., 2025), these methods were not specifically designed for supporting senior dialogue engagement. Our approach is tuning-free yet compatible with RL-driven methods, enabling future extensions through hierarchical RL to further enhance multi-turn conversation engagement.
3 ChatWise Design
ChatWise employs a dual-level framework with a strategy provider $\pi_{s}$ that induces macro-level actions, and an utterance generator $\pi_{u}$ for deriving utterances based on suggested strategies (Figure 1).
Strategy Candidates: To construct a strategy candidate set, we extract Dialogue Acts (DAs) from real telehealth clinical trials Yuan et al. (2023), where DAs serve as atomic communicative units that convey distinct conversational intentions Searle (1976). We further integrate these with strategies from prior emotional support dataset (Liu et al., 2021a) to form a comprehensive set. We refine each strategy with definitions and examples as in-context prompts to enhance $\pi_{s}$ reasoning.
From Strategy to Utterance: $\pi_{s}$ reasons over the dialogue history $\tau$ to capture user’s emotion $e$ and derive one to two strategies, $\{s_{i}\}_{i\leq 2}\sim\pi_{s}(\tau,e)$. Our rationale follows counseling study (Zhang and Danescu-Niculescu-Mizil, 2020) that dialogue intentions can be forward, e.g. initiating new topics via an open question, or backward, e.g. responding with acknowledgment. $\pi_{s}$ is prompted to flexibly employ one or both as needed. Selected strategies, user’s current emotion, and conversation history serve as inputs for $\pi_{u}$ to generate utterances. To ensure natural flow, $\pi_{u}$ will first improvise a few rounds before following suggested strategies. Drawing on clinical guidance, $\pi_{u}$ encourages users to choose topics rather than imposing them.
4 Empirical Evaluation
Below we summarize experiment configurations, metrics, and main findings of ChatWise. More experimental details are available in Appendix A.1 and A.2.
4.1 Conversation Generation
For controlled and responsive development of ChatWise, we employed 9 digital twins provided by  Hong et al. (2024) as simulated users with different personas, which are fine-tuned LLMs to approximate senior individuals with MCI symptoms using real, de-identified dialogue data from a clinical trial  I-CONECT (2024). We collected 20 trajectories of conversations between ChatWise and each digital twin, where a conversation contains 10 turns, each turn with a two-way utterance exchange between ChatWise and digital twin (user).
4.2 Engagement Evaluation
User verbosity: We primarily measure user engagement through verbosity, i.e. the user-to-chatbot token ratio, defined as $\frac{\text{tokens}_{user}}{\text{tokens}_{chatbot}}$, as a clinical study (Yu et al., 2021) suggested reducing the moderator talkativeness while encouraging participant (user) expression.
Cognitive Metrics: See et al. (2019) defines 9 aspects for evaluating conversation quality, from which we select 5 focused on engagement assessment: Engagingness, Interestingness, Listening, Fluency, and Making Sense. While Interestingness assesses overall dialogue quality, the others primarily evaluate user interactions.
We compute the win rate (Rafailov et al., 2023) by comparing dialogues generated with and without ChatWise, using the same pretrained LLM (GPT-4o) as the utterance generator and identical system prompts, except for input from $\pi_{s}$ to $\pi_{u}$. Dialogue pairs are randomly matched within the same digital twin for evaluation.
Strategy Provider Verbosity $\uparrow$ Engagingness WR $\uparrow$ Interestingness WR $\uparrow$ Listening WR $\uparrow$ Fluency WR $\uparrow$ Making sense WR $\uparrow$
w/o ChatWise 0.7398 0.3389 0.3833 0.3593 0.4111 0.4056
GPT-4o 0.8635 0.6833 0.6222 0.6222 0.5778 0.5833
o3-mini 0.8643 0.6778 0.6222 0.6778 0.6222 0.6167
Llama 3.1-405B 0.8083 0.6222 0.6056 0.6222 0.5667 0.5833
Table 1: Performance on 10-turn dialogue across different strategy providers. W/o ChatWise is the baseline without applying any strategy, whose win rates (WRs) are averaged across strategy providers. (Appendix A.3)
4.3 ChatWise Performance and Analysis
We evaluated ChatWise using different LLM backbones as strategy providers, including GPT-4o, o3-mini, and Llama3.1-405B. As shown in Table 1, ChatWise consistently outperformed non-strategic dialogues across all evaluation metrics.
ChatWise robustly enhances user engagement: Regardless of the capability of pretrained LLMs, our design consistently improves user engagement and dialogue quality with a large margin across all tested LLMs as the strategy provider.
Reasoning ability can transfer to improve engagement: The best-performing model in our setting, o3-mini, was optimized for STEM reasoning tasks, which indicates that strong reasoning ability can also benefit inferring dialogue strategies.
4.4 Accumulated Multi-Turn Analysis
We analyzed ChatWise’s performance over increasing conversation turns, as shown in Figure 3, which presents accumulated metrics and win rates. Results revealed a clear upward trend in all key engagement metrics—especially verbosity, engagingness, interestingness, and listening—demonstrating that ChatWise’s effects strengthen as dialogues progress, while baseline gradually degrades, which highlights ChatWise’s capability in enhancing long-term conversational engagement.
(a) Verbosity.
(b) Engagingness Win Rate.
(c) Interestingness Win Rate.
(d) Listening Win Rate.
Figure 3: ChatWise performance with o3-mini as strategy provider. W/o ChatWise is the baseline and GPT-4o servers as utterance generator in all settings. Evaluation starts by skipping two warming-up turns (Sec 3).
4.5 Ablation Study
We evaluated ChatWise without user emotion in the strategy provider’s output (Table 2). Its consistent performance confirms that strategies are the primary driver of engaging conversations.
Strategy Provider GPT-4o GPT-4o w/o emotion w/o ChatWise
Verbosity 0.8635 0.8463 0.7398
Table 2: User verbosity under different strategy providers. GPT-4o w/o emotion keeps the strategy provider with user emotion information removed.
4.6 ChatWise’s Robustness to User Persona
Figure 4 shows the significant gain of ChatWise interacting with all simulated users, indicating its robustness to support varying senior characteristics.
Figure 4: Users’ dialogue verbosity (log-normalized, Appendix A.5) across 9 digital twins using different LLMs as strategy providers, compared with baseline (w/o ChatWise).
4.7 Dialogue Analysis
We identified 4 digital twins for which ChatWise was most effective and collected 40 additional dialogues from each to conduct a deeper analysis, with the main findings summarized below.
Transient user emotions: We defined emotion transition triplets as $(\text{e}_{i},\text{s},\text{e}_{i+1})$, where $\text{e}_{i}$ ($\text{e}_{i+1}$) represents the user’s emotion at turn $i$ ($i+1$), and $s$ is the strategy provided by ChatWise at turn $i$. We computed the average occurrence of emotion triplets and showed the top 15 most frequent in Figure 5. Most triplets reflect unchanged user emotions, indicating the challenge of detecting or influencing emotional shifts within a short turn.
Long-term user emotions: We analyzed user emotion changes from the beginning to the end of dialogues and calculated averaged occurrence across digital twins (Figure 6). Over 48% users experienced emotional shifts post-dialogue, with significantly more positive changes after engaging with ChatWise. This implies that while transient emotions are difficult to track, strategic conversational support can improve user emotions over time.
Predominant strategies: We identified the top 10 most frequent strategies across digital twins and found that predominant interaction strategies remained consistent across user types (See Appendix A.8). Particularly, Open Question, Statement-non-opinion, and Acknowledgment strategies dominate ChatWise driven conversations, suggesting their potential efficacy in fostering engaging dialogues.
Figure 5: Average occurrence of each emotion transition triplet across the samples of each digital twin.
Figure 6: Occurrence of user emotional changes. For instance, “neutral-joy” indicates user emotion transitioning from neutral (conversation begins) to joy (ends).
5 Conclusion and Future Work
We introduce ChatWise, a tuning-free yet principled framework for developing AI chatbots that foster long-term user engagement. Our dual-level generation design integrates clinical insights into LLM reasoning and is compatible with various LLMs. Extensive evaluations based on de-identified clinical data showed that our method robustly enhances the engagement and cognitive status of seniors. Moving forward, we aim to 1) enhance multi-turn interaction via RL and 2) validate its effectiveness through real user studies.
Limitations
All digital twins were provided as fine-tuned GPT-3.5 APIs, and its high economic cost, at the time when this project was going, constrained the volume of dialogues we generated. To mitigate this, future work will explore training digital twins using LLaMA 3.1 to reduce sampling costs while maintaining conversational quality. Additionally, our study relies on simulated dialogues rather than real user interactions. A real user study is needed to validate the system’s effectiveness in dynamic settings, which we plan to incorporate into our future research.
Appendix A Appendix
A.1 Prompt design
The prompt used in the experiments and structured output design are available at https://anonymous.4open.science/r/ChatWise-D5E0, including:
  • Strategy provider system prompt.
  • Moderator initial system prompt.
  • Moderator system prompt with strategies.
  • Strategy provider system prompt for ablation study.
  • Moderator system prompt with strategies for ablation study.
  • Structured output class for OpenAI models as strategy provider.
  • Structured output class for OpenAI models as strategy provider in ablation study.
  • System prompt for GPT-4o to extract the strategy given by Llama3.1.
A.2 Data Generation Configuration
We used GPT-4o as the utterance generator in ChatWise and tested different LLM backbones as strategy providers. Considering the Llama3.1-405B does not support structured output, we applied GPT-4o as the strategy extractor to structure the strategy output of Llama3.1-405B. The following are the configurations of each model:
Utterance generator (GPT-4o):
n=1
max_tokens=1024
top_p=1
temperature=1
GPT-4o, o3-mini as strategy provider:
n=1
max_tokens=1024
top_p=1
temperature=1
response_format=Strategy
Llama3.1-405B as strategy provider:
top_p: 0.9
max_tokens: 1024
temperature: 0.6
presence_penalty: 0
frequency_penalty: 0
GPT-4o as strategy extractor:
n=1
max_tokens=1024
top_p=1
temperature=1
A.3 Win rate for w/o ChatWise
The Win rate for w/o ChatWise is defined as:
$\begin{split}1-\text{average}(&\text{WR}_{GPT-4o}\\ &+\text{WR}_{o3-mini}\\ &+\text{WR}_{Llama3.1-405B})\end{split}$
,where $\text{WR}_{X}$ is ChatWise’s Win rate against the baseline with $X$ as strategy provider.
A.4 Dialogue Acts
The map of strategy to its corresponding abbreviated tag is listed in Table3.
Strategy Tag
Acknowledge (Backchannel) Ack
Statement-non-opinion StaNo
Statement-opinion Sta
Affirmation and Reassurance Agr
Appreciation App
Conventional-closing ConC
Hedge H
Other Oth
Quotation Quo
Action-directive AcD
Collaborative Completion CoC
Restatement or Paraphrasing Rep
Offers Options Commits Off
Self-talk Sel
Apology Apo
Reflection of Feelings RoF
Yes-No-Question YNQ
Wh-Question WhQ
Declarative Yes-No-Question DYNQ
Open-Question OpQ
Or-Clause OrC
Conventional-opening CoO
Self-disclosure Sd
Providing Suggestions PS
Information I
Table 3: Dialogue Act to its corresponding tag.
There are two kinds of strategies: backward-looking and forward-looking. Backward-looking strategies reflect how the current utterance relates to the previous discourse. Forward-looking strategies reflect the current utterance constrains the future beliefs and actions of the participants and affects the discourse. Table 4 and Table 5 provide definitions and examples of each.
Strategy Definiton Example
StaNo A factual statement or descriptive utterance that does not include an opinion. Me, I’m in the legal department.
Ack A brief utterance that signals understanding, agreement, or active listening. Uh-huh.
Sta A statement that conveys a personal belief, judgment, or opinion. I think it’s great
Agr Affirm the help seeker’s strengths, motivation, and capabilities and provide reassurance and encouragement. That’s exactly it.
App An expression of gratitude, admiration, or acknowledgment of another’s effort or input. I can imagine.
ConC A formal or socially standard utterance signaling the end of a conversation. Well, it’s been nice talking to you.
H An expression that introduces uncertainty or qualification to a statement, often to soften its impact. I don’t know if I’m making any sense or not.
Oth Exchange pleasantries and use other support strategies that do not fall into the above categories. Well give me a break, you know.
Quo A direct or indirect repetition of someone else’s words. Albert Einstein once said, “Imagination is more important than knowledge.”
AcD A command, request, or suggestion directing someone to take action. Why don’t you go first
CoC A continuation or completion of someone else’s utterance in a collaborative manner. If we want to make it to the top of the mountain before sunset, we should…
Rep A simple, more concise rephrasing of the help-seeker’s statements that could help them see their situation more clearly. It sounds like you’re saying that you’re struggling to stay on top of your work, and it’s leaving you feeling overwhelmed.
Off A statement proposing choices, making a commitment, or offering to do something. I’ll have to check that out
Sel An utterance directed at oneself, often reflecting internal thought processes or problem-solving. What’s the word I’m looking for
Apo An expression of regret or asking for forgiveness. I’m sorry.
RoF Articulate and describe the help-seeker’s feelings. It sounds like you’re feeling really frustrated and drained because your efforts don’t seem to be paying off.
Table 4: Backward-looking strategies, definition, and example.
Strategy Definiton Example
YNQ A question expecting a binary (yes/no) response. Do you have to have any special training?
WhQ A question beginning with a wh-word (e.g., what, who, where), seeking specific information. Well, how old are you?
DYNQ A statement posed as a question, expecting a yes/no answer. So you can afford to get a house?
OpQ A broad question inviting a wide range of responses, often conversational. How about you?
OrC A question offering explicit alternatives, often in the form of “or.” or is it more of a company?
CoO A socially standard utterance used to initiate a conversation. How are you?
Sd Divulge similar experiences that you have had or emotions that you share with the help-seeker to express your empathy. I completely understand how you feel. I remember feeling the same way before my first big presentation at work. I was so anxious, but I found that practicing a few extra times really helped calm my nerves.
PS Provide suggestions about how to change, but be careful to not overstep and tell them what to do. You can keep a note to stop your idea from going.
I Provide useful information to the help-seeker, for example with data, facts, opinions, resources, or by answering questions. Taking silver line from Washington D.C. to Dulles Intel Airport costs about 1 hour.
Table 5: Forward-looking strategies, definition, and example.
A.5 Log-normalization
The following is the log-normalization function, where y is the normalized result, x is the input variable.
$y=\ln{\left(4x+1\right)}$
A.6 Additional results for Accumulated Multi-Turn Analysis
ChatWise’s accumulated performance on Fluency, and Making Sense is shown in Figure 7.
(a) Fluency Across Turns
(b) Making sense Across Turns
Figure 7: ChatWise performance with o3-mini as strategy provider across turns. W/o ChatWise is the baseline that ChatWise is not applied. GPT-4o servers as moderator in all settings.
A.7 Primary Strategies
We calculated the average occurrence of each strategy across each digital twin and listed their top 10 most frequently occurring strategies, as shown in Figure 8.
Figure 8: Strategy occurrence across digital twins.
A.8 Personalized Dialogue Analysis
This section shows the personalized dialogue analysis. The Strategy occurrence across digital twins is shown in Figure 9. The Occurrence of each emotion transition triplet across digital twins is shown in Figure 10. The Open Question, Statement-non-opinion, and Acknowledgment strategies still dominate ChatWise driven conversations, suggesting their potential effectiveness in fostering engagement in conversations. The user’s emotion is detected as unchanged in most triplets, indicating the difficulty of altering or measuring the user’s emotional movement within a short turn.
(a) Digital Twin 3.
(b) Digital Twin 5.
(c) Digital Twin 6.
(d) Digital Twin 9.
Figure 9: Strategy occurrence across digital twins.
(a) Digital Twin 3.
(b) Digital Twin 5.
(c) Digital Twin 6.
(d) Digital Twin 9.
Figure 10: Occurrence of each emotion transition triplet across digital twins.

### v2

@@
-ChatWise: AI-Powered Engaging Conversations for Enhancing Senior Cognitive Wellbeing
+ChatWise: A Strategy-Guided Chatbot for Enhancing Cognitive Support in Older Adults
 Abstract
 Cognitive health in older adults presents a growing challenge.
-While conversational interventions show feasibility in improving cognitive wellness, human caregiver resources remain overburdened.
-AI-based methods have shown promise in providing conversational support, yet existing work is limited to implicit strategy while lacking multi-turn support tailored to seniors.
-We improve prior art with an LLM-driven chatbot named ChatWise  for older adults.
-It follows dual-level conversation reasoning at the inference phase to provide engaging companionship.
-ChatWise thrives in long-turn conversations, in contrast to conventional LLMs that primarily excel in short-turn exchanges.
-Grounded experiments show that ChatWise significantly enhances simulated users’ cognitive and emotional status, including those with Mild Cognitive Impairment.
-ChatWise: AI-Powered Engaging Conversations
-for Enhancing Senior Cognitive Wellbeing
-Zhengbang Yang, Zhuangdi Zhu George Mason University, Fairfax, VA, USA {zyang30,zzhu24}@gmu.edu
+Although conversational interventions show feasibility in improving cognitive wellness, human caregiver resources remain overloaded.
+AI-based chatbots have shown promise, yet existing work is often limited to implicit strategies or heavily depends on training and label resources.
+In response, we propose a strategy-guided AI chatbot named ChatWise  that follows a dual-level conversation reasoning framework.
+It integrates macro-level strategy planning and micro-level utterance generation to enable engaging, multi-turn dialogue tailored to older adults.
+Empirical results show that ChatWise closely aligns with professional human caregiver behaviors in offline evaluation using real clinic data, and achieves positive user cognitive and emotional responses in interactive simulations with digital twins, which significantly outperforms AI baselines that follow implicit conversation generation.
+ChatWise: A Strategy-Guided Chatbot for Enhancing Cognitive Support in Older Adults
+Zhengbang Yang1, Junyuan Hong2, Yijiang Pang3, Jiayu Zhou4, Zhuangdi Zhu1 1 George Mason University, Fairfax, VA, USA 2 University of Texas at Austin, Austin, TX, USA 3 Michigan State University, East Lansing, MI, USA 4 University of Michigan, Ann Arbor, MI, USA {zyang30,zzhu24}@gmu.edu
 1 Introduction
-[…]
-In this work, we aim to draw on the advance of AI to provide older adults with engaging conversational support that improves their cognitive function and reduces social loneliness, serving as accessible alternatives to human companions.
-Previous clinical studies aimed at socially isolated older adults showed the existence of causal relationships between interviewer strategy and interviewee response  (Cao et al., 2021), revealing that conversation behavior can have a measurable influence on interlocutors, which inspired us to develop principled yet efficient methods that transform clinical insights into strategic dialogue design leveraging the advanced reasoning ability of LLMs.
+The cognitive well-being of the elderly population is a pressing social concern, as evidenced by the prevalence of cognitive disorders within this population, often exacerbated by loneliness and isolation Nicholson (2012); Teo et al.
+(2023).
+According to WHO, approximately $14\%$ of adults aged 60 and over experience mental health disorders, projected to affect 2.1 billion individuals by 2050 Organization (2024).
+Compared with other age groups, older adults are more vulnerable due to age-related changes in cognitive reserves  (Salthouse, 2009) and reduced social connections  (Nicholson, 2012; Teo et al., 2023).
+Such impact extends beyond older individuals to families and society, resulting in a reduced life quality and increased medical burden.
+In the meantime, modest delays in cognitive decline can significantly reduce dementia prevalence, and addressing social isolation could prevent 4% of dementia cases Livingston et al.
+(2020).
+Interventions through guided conversations have shown efficacy in reducing loneliness and mitigating cognitive decline  (Yu et al., 2021, 2022, 2023; Fiori and Jager, 2012).
+However, effective intervention requires sustained interaction and monitoring, which is limited by the availability of human companions, leading to inconsistent access or effectiveness.
+Advances in artificial intelligence (AI) particularly Large Language Models (LLMs) have shown promise in augmenting human expertise with conversational support.
+Existing efforts range from sentiment-based chatbots Liu et al.
+(2021); sen (2024); Ryu et al.
+(2020); Tsai et al.
+(2021); You et al.
+(2023); Meyer et al.
+(2020) to audio assistants Yang et al.
+(2024); Hong et al.
+(2024), ascribed to mature text-audio transformation techniques Van Den Oord et al.
+(2016); Ren et al.
+(2019).
+However, these systems primarily default to interactions with implicit goals, which may fail to drive engaging conversations that necessitate strategic multi-turn interactions tailored to older adults.
+Some work has explored fine-tuning LLMs on domain-specific datasets to adapt to elderly care Sun et al.
+(2025), which requires extensive labeled data and resources, while they may struggle to generalize across conversational contexts.
+In contrast, we take an orthogonal approach by focusing on the inherent reasoning capabilities of LLMs through in-context learning, which offers more resource efficiency and adaptability.
+Our goal is to provide AI-powered, engaging conversational support for older adults that serves as an accessible complement to human companions, with the aspiration to improve their cognitive function and reduce social loneliness.
+Previous clinical studies aimed at socially isolated older adults showed the existence of causal relationships between interviewer strategy and interviewee response  (Cao et al., 2021), which revealed that conversation behavior can have a measurable influence on interlocutors, and thus inspired us to develop principled yet efficient methods that transform clinical insights into dialogue design by leveraging the advanced reasoning ability of LLMs.
 In response, we propose an LLM-driven chatbot named ChatWise.
-Unlike traditional chatbot driven with implicit conversational goals, ChatWise employs dual-layer conversation generation, which first derives categorized macro-level information, including user emotion states to suggest meta-conversational strategies, which then guides the micro-level utterance generation to improve both user engagement and cognitive outcomes over multi-turn dialogue interactions.
-Figure 1 overviews its development design.
-Our work provides multifold contributions:
-  • ChatWise is a tuning-free yet principled framework for developing AI chatbots with enhanced multi-turn, daily-themed conversational support.
-Our dual-level conversation generation design integrates clinical insights into LLM reasoning and can be readily applied to various LLMs.
-  • ChatWise follows grounded development and evaluation using digital twins (Hong et al., 2024), which are simulated users constructed by tuning LLMs on de-identified, real-world dialogues between seniors and professional caregivers of a clinical trial (I-CONECT, 2024).
-Interactions with digital twins demonstrated that ChatWise significantly enhances users’ engagement and cognitive status with multi-turn interactions, especially for users with Mild Cognitive Impairment (MCI) (Figure 2).
-  • Our comparative studies demonstrated that integrating macro-level strategies into conversation generation is the key contributor to enhancing user engagement.
-Its gain on cognitive metrics remains consistent across different backbone LLMs.
-ChatWise’s advantages become more pronounced in longer conversations, in contrast to conventional LLMs that primarily excel in short-turn question-answer exchanges.
-  • Our dialogue analysis revealed key empirical insights, including the pattern of predominant interaction strategies across users.
-While transient user emotional fluctuations were less tractable, multi-turn conversations with ChatWise have shown to notably improve user emotions over time.
-This highlights the value of long-term, principled conversational support in enhancing seniors’ cognitive and emotional well-being.
-We hope these findings can inspire future research, including RL-based LLM optimization.
-Figure 1: Overview of ChatWise Development.
-Figure 2: ChatWise notably improves multiple user cognitive metrics through 10-turn dialogues (Sec 4.2).
+It employs dual-layer conversation generation that first derives categorized macro-level information to suggest meta-conversational strategies, which then guides the micro-level utterance generation to improve both user engagement and cognitive outcomes over multi-turn dialogue interactions.
+ChatWise is evaluated using both real, de-identified real-world dialogues between older adults and professional human caregivers of a clinical trial (I-CONECT, 2024), and synthetic interactive conversations using simulated users (i.e.
+digital twins)  (Hong et al., 2024) modeled on such data.
+Our work provides multifold contributions: (i) We introduce a clinically grounded chatbot that highly aligns with the behavior of professional human caregivers to older adults (Sec 4.1).
+Figure 1 overviews its alignment performance.
+It also empirically enhances simulated users’ engagement and cognitive status, which significantly outperforms baseline chatbots (Sec 4.2).
+(ii) ChatWise follows a tuning-free, in-context learning framework for daily conversational support.
+Comparative studies demonstrated that providing macro-level strategies to guide conversation generation is the key contributor to enhancing user engagement.
+(iii) Its dual-level policy design can be readily applied to various LLMs, with consistent user cognitive gains across different backbone LLMs.
+Figure 1: Alignment between conversational strategies of AI chatbots and human professionals across diverse participant data.
+Our design shows closer alignment with caregiver behavior in real conversation contexts compared to baseline chatbots.
+See Sec 4 for full results.
 2 Related Work
-AI-powered Chatbots for Seniors: Social interaction plays a crucial role in the cognitive health of seniors  (Patil and Braun, 2024; Fiori and Jager, 2012).
-AI-powered chatbots have emerged as a promising alternative  (Rodríguez-Martínez et al., 2023; Owan et al., 2023).
-Recent efforts span commercial products (sen, 2024; ell, 2024) and research prototypes focused on emotional support and audio assistance  (Yang et al., 2024; Liu et al., 2021b; sen, 2024; Ryu et al., 2020; Hong et al., 2024).
-Unlike prior work, our approach prioritizes user conversational engagement, which has been empirically shown to enhance both cognitive and emotional status.
-Conversational Strategies: Liu et al.
-(2021a) curated a dataset with annotated strategies, demonstrating the effectiveness of Helping Skills Theory Hill (2020) in providing emotional support.
+AI-powered Chatbots for Older Adults
+have shown feasibility in preventing or detecting the cognitive issues to assist human professionals.
+Recent efforts span commercial products (sen, 2024; ell, 2024) and research prototypes focused on emotional support and audio assistance.
+(Sun et al., 2025) performed supervised fine-tuning on LLMs to enhance their performance on specialized nursing and elderly care tasks.
+Yang et al.
+(2024) introduced an LLM-based voice assistant designed to bridge communication between older adults and their healthcare providers.
+Liu et al.
+(2021) built an annotated dataset to tune LLMs for emotional support tasks.
+Ryu et al.
+(2020) developed a chatbot for the mental health of the elderly.
+Hong et al.
+(2024) constructed digital twins of the elderly.
+Unlike prior work, our approach prioritizes user conversational engagement through in-context reasoning, which has been empirically shown to enhance user cognitive status.
+Dialogue Systems for Mental Health or Cognitive Stimulation:
+Recent studies have leveraged LLMs for the augmentation of emotional support conversations through generated dialogue data  (Zheng et al., 2023; Jiang et al., 2023).
+Xygkou et al.
+(2024) assessed the acceptability of GPT-4-based conversational agents among people with dementia (PwD), and highlighted the importance of design considerations for this sensitive population.
+Similarly,  Favela et al.
+(2023) implemented a conversational robot embedded with ChatGPTs for reminiscence therapy, with authors as simulated users, which showcased the potential of generative AI to support cognitive stimulation through interactive storytelling.
+Distinct from these works, our research introduces a dual-level dialogue approach combined with comprehensive evaluation methodologies to provide a finer-grained understanding of conversational engagement.
+Conversational Strategies:
+Liu et al.
+(2021) curated a dataset with annotated strategies, demonstrating the effectiveness of Helping Skills Theory Hill (2020) in providing emotional support.
 Yuan et al.
@@
 (2021) identified key strategies for improving child patient-provider communication through semi-structured interviews.
-Few works have systematically integrated these strategies into the reasoning flow of AI chatbots.
-Our approach fills this gap by embedding structured conversational strategies for automatically enhancing user engagement.
-Multi-turn chatbot exploration: Recent advances in AI dialogue systems predominantly focused on short-turn interactions  (Owan et al., 2023; Dam et al., 2024).
+However, few works have systematically integrated these strategies into the reasoning flow of AI chatbots.
+Our approach fills this gap by contextualizing structured conversational strategies for automatically enhancing user engagement.
+Multi-Turn Chatbot Exploration:
+Recent advances in AI dialogue systems predominantly focused on short-turn interactions  (Owan et al., 2023; Dam et al., 2024; Gao et al., 2025), with relatively limited attention to the challenges of multi-turn exchanges.
 While some pioneering research has explored multi-turn optimization through Reinforcement Learning (RL) approaches for LLMs (Verma et al., 2022; Zhou et al., 2024; Abdulhai et al., 2023; Gao et al., 2025), these methods were not specifically designed for supporting senior dialogue engagement.
-Our approach is tuning-free yet compatible with RL-driven methods, enabling future extensions through hierarchical RL to further enhance multi-turn conversation engagement.
+Our inference-based method introduces a strategy-compatible framework that aligns with the principles of RL-driven optimization, which can enable future extension to further enhance multi-turn conversation engagement.
+Figure 2: ChatWise  employs a dual-level policy design that contextualizes a clinically-derived strategy pool to generate macro-level actions, which then guide micro-level utterance generation.
+It is iteratively developed using digital twins as simulated users.
+Prior to a future real user study, real and simulated data analysis show that ChatWise closely aligns with professional human caregivers’ behavior, leading to cognitive gains of simulated users accumulated over multi-turn interactions.
 3 ChatWise Design
-ChatWise employs a dual-level framework with a strategy provider $\pi_{s}$ that induces macro-level actions, and an utterance generator $\pi_{u}$ for deriving utterances based on suggested strategies (Figure 1).
-Strategy Candidates: To construct a strategy candidate set, we extract Dialogue Acts (DAs) from real telehealth clinical trials Yuan et al.
-(2023), where DAs serve as atomic communicative units that convey distinct conversational intentions Searle (1976).
-We further integrate these with strategies from prior emotional support dataset (Liu et al., 2021a) to form a comprehensive set.
-We refine each strategy with definitions and examples as in-context prompts to enhance $\pi_{s}$ reasoning.
-From Strategy to Utterance: $\pi_{s}$ reasons over the dialogue history $\tau$ to capture user’s emotion $e$ and derive one to two strategies, $\{s_{i}\}_{i\leq 2}\sim\pi_{s}(\tau,e)$.
-Our rationale follows counseling study (Zhang and Danescu-Niculescu-Mizil, 2020) that dialogue intentions can be forward, e.g.
+ChatWise encompasses a hierarchical framework that integrates the insights of traditional clinical studies to drive LLM reasoning.
+It features a dual-level design with a strategy provider $\pi_{s}$ that generates macro-level conversation strategies, which guides a utterance generator $\pi_{u}$ to produce micro-level responses.
+In this paper, we interchangeably refer to ChatWise as the interviewer or the moderator, a role commonly defined in communication studies that guides and facilitates the conversation  (Taboada, 2006),
+For clarity, we denote a multi-turn conversation as $\tau=\{\mu_{0},u_{0},x_{1},u_{1},x_{2},u_{2},\cdots\}$ which comprises interviewee (user) utterances $x_{t}$ and interviewer (ChatWise) utterances $u_{t}$ at time step $t$, with a conversation starting point $\mu_{0}$.
+$s_{t}=\{\mu_{0},u_{0},\cdots,x_{t}\}$ denotes the historical conversation up to timestep $t$.
+We then formulate ChatWise as a dual policy $\pi=\{\pi_{s},\pi_{u}\}$, which maps the conversational state $s_{t}$ to the next utterance $u_{t}$ provided to the user, conditioned on meta information $z$ given to the strategy provider $\pi_{s}$:
+$u_{t}\sim\pi(\cdot|s_{t},z)\equiv\pi_{u}(\cdot|\pi_{s}(s_{t}|z)).$
+Below, we illustrate the process of this hierarchical conversation generation.
+3.1 Strategy Pool
+To first develop clinically grounded strategies, we establish a strategy pool containing dialogue acts (DA) curated from clinical intervention studies involving older adults.
+These DAs, such as asking an open question (e.g.
+“What is your favorite movie”) or showing empathy (e.g.“That sounds really meaningful”), serve as macro-level actions and atomic communicative units that convey distinct conversational intentions Searle (1976), which have also shown causal relationships with participant emotions in conversations Cao et al.
+(2021).
+Specifically, we extract DAs that exist in conversations from a real tele-health clinical trial Yuan et al.
+(2023), which represents a subset of the complete DAMSL taxonomy Allen and Core (1997); Searle (1976).
+We further augment this DA with strategies from prior emotional support dialogue dataset (Liu et al., 2021) to form a comprehensive strategy pool $\mathcal{A}$.
+We then treat each strategy serves as a macro-level action candidate $a\in\mathcal{A}$ and refine it with a definition and an example to serve as in-context information $z$ for enhancing the reasoning of strategy $\pi_{s}$ (See Appendix A.5 for a complete DA set).
+3.2 Emotion Annotation
+To enhance context-awareness during conversations and support post-conversation analysis, the strategy provider $\pi_{s}$ is additionally prompted to infer the user’s current emotion based on historical conversations $s_{t}$.
+At timestep $t$ of the conversation, $\pi_{s}$ will categorize user emotion into one of the five classes: joy, neutral, sadness, anger, and surprise, given a calibrated system prompt.
+Our comparative study shows that ChatWise can lead to detectable positive emotion changes over multi-turn conversations, which aligns with conversations provided by human professionals (Sec 4.5).
+3.3 From Strategy to Utterance Generation
+The strategy provider $\pi_{s}$ processes the dialogue history $s_{t}$, contextualized by the strategy pool information $z$ to decide the most appropriate strategies for continuing a conversation.
+We limit each interviewer’s utterance to contain one or two strategies, i.e.
+$\mathbf{a}_{t}=\{a_{t}^{i}\}_{i\leq 2}\sim\pi_{s}(\cdot|s_{t},z)$.
+Our rationale follows counseling studies (Zhang and Danescu-Niculescu-Mizil, 2020) that dialogue intentions can be forward, e.g.
 initiating new topics via an open question, or backward, e.g.
-responding with acknowledgment.
-$\pi_{s}$ is prompted to flexibly employ one or both as needed.
-Selected strategies, user’s current emotion, and conversation history serve as inputs for $\pi_{u}$ to generate utterances.
-To ensure natural flow, $\pi_{u}$ will first improvise a few rounds before following suggested strategies.
-Drawing on clinical guidance, $\pi_{u}$ encourages users to choose topics rather than imposing them.
-4 Empirical Evaluation
-Below we summarize experiment configurations, metrics, and main findings of ChatWise. More experimental details are available in Appendix A.1 and A.2.
-4.1 Conversation Generation
-For controlled and responsive development of ChatWise, we employed 9 digital twins provided by  Hong et al.
-(2024) as simulated users with different personas, which are fine-tuned LLMs to approximate senior individuals with MCI symptoms using real, de-identified dialogue data from a clinical trial  I-CONECT (2024).
-We collected 20 trajectories of conversations between ChatWise and each digital twin, where a conversation contains 10 turns, each turn with a two-way utterance exchange between ChatWise and digital twin (user).
-4.2 Engagement Evaluation
-User verbosity: We primarily measure user engagement through verbosity, i.e.
-the user-to-chatbot token ratio, defined as $\frac{\text{tokens}_{user}}{\text{tokens}_{chatbot}}$, as a clinical study (Yu et al., 2021) suggested reducing the moderator talkativeness while encouraging participant (user) expression.
-Cognitive Metrics: See et al.
-(2019) defines 9 aspects for evaluating conversation quality, from which we select 5 focused on engagement assessment: Engagingness, Interestingness, Listening, Fluency, and Making Sense.
-While Interestingness assesses overall dialogue quality, the others primarily evaluate user interactions.
-We compute the win rate (Rafailov et al., 2023) by comparing dialogues generated with and without ChatWise, using the same pretrained LLM (GPT-4o) as the utterance generator and identical system prompts, except for input from $\pi_{s}$ to $\pi_{u}$.
-Dialogue pairs are randomly matched within the same digital twin for evaluation.
-Strategy Provider Verbosity $\uparrow$ Engagingness WR $\uparrow$ Interestingness WR $\uparrow$ Listening WR $\uparrow$ Fluency WR $\uparrow$ Making sense WR $\uparrow$
-w/o ChatWise 0.7398 0.3389 0.3833 0.3593 0.4111 0.4056
-GPT-4o 0.8635 0.6833 0.6222 0.6222 0.5778 0.5833
-o3-mini 0.8643 0.6778 0.6222 0.6778 0.6222 0.6167
-Llama 3.1-405B 0.8083 0.6222 0.6056 0.6222 0.5667 0.5833
-Table 1: Performance on 10-turn dialogue across different strategy providers.
-W/o ChatWise is the baseline without applying any strategy, whose win rates (WRs) are averaged across strategy providers.
-(Appendix A.3)
-4.3 ChatWise Performance and Analysis
-We evaluated ChatWise using different LLM backbones as strategy providers, including GPT-4o, o3-mini, and Llama3.1-405B.
-As shown in Table 1, ChatWise consistently outperformed non-strategic dialogues across all evaluation metrics.
-ChatWise robustly enhances user engagement: Regardless of the capability of pretrained LLMs, our design consistently improves user engagement and dialogue quality with a large margin across all tested LLMs as the strategy provider.
-Reasoning ability can transfer to improve engagement: The best-performing model in our setting, o3-mini, was optimized for STEM reasoning tasks, which indicates that strong reasoning ability can also benefit inferring dialogue strategies.
-4.4 Accumulated Multi-Turn Analysis
-We analyzed ChatWise’s performance over increasing conversation turns, as shown in Figure 3, which presents accumulated metrics and win rates.
-Results revealed a clear upward trend in all key engagement metrics—especially verbosity, engagingness, interestingness, and listening—demonstrating that ChatWise’s effects strengthen as dialogues progress, while baseline gradually degrades, which highlights ChatWise’s capability in enhancing long-term conversational engagement.
-(a) Verbosity.
-(b) Engagingness Win Rate.
-(c) Interestingness Win Rate.
-(d) Listening Win Rate.
-Figure 3: ChatWise performance with o3-mini as strategy provider.
-W/o ChatWise is the baseline and GPT-4o servers as utterance generator in all settings.
-Evaluation starts by skipping two warming-up turns (Sec 3).
+responding with acknowledgment, or both to transit between intentions or topics.
+Thus, $\pi_{s}$ is prompted to comply with one of the three conditions: (i) a forward strategy, (ii) a backward strategy, or (iii) a backward strategy followed by a forward strategy.
+The selected strategies $\mathbf{a}_{t}$, and optionally, the user’s current emotion label $e_{t}$, and the conversation history $s_{t}$ are used as inputs to the utterance generator $\pi_{u}$.
+Our ablation study showed that providing meta-level strategies has a dominant effect over providing emotion information alone for conversational engagement (Sec 4.3).
+To ensure natural flow, $\pi_{u}$ will first improvise a few rounds as warm-ups before adopting the suggested strategies.
+We also draw on clinical guidance to let $\pi_{u}$ encourage users to choose topics rather than imposing them.
+Figure 2 overviews this design.
+4 Experiments
+We conducted various experiments focusing on answering the following questions:
+Q1. Does ChatWise’s behavior align with human professionals when responding to real conversation context sampled from clinic trials?
+Q2. How does ChatWise compare to baseline AI alternatives in terms of multifaceted conversational engagement metrics?
+Q3. How do factors such as the conversation turns and different user characteristics influence ChatWise’s performance?
+To address these questions, we evaluated ChatWise on two complementary scenarios: we first conducted offline evaluations using data from a real clinical trial (Sec 4.1), then we performed interactive experiments, in which conversations are stochastically generated between ChatWise and digital twins as simulated users (Sec 4.2).
+Below, we summarize the configurations, metrics, and main results of each experimental setting and defer more details to the Appendix.
+4.1 Offline Evaluation on Real Clinical Data
+4.1.1 Data Preprocessing
+We extracted dialogues from the I-CONECT clinic trials I-CONECT (2024), with each in text format converted from video conversations.
+Short dialogues below 40 turns were excluded, where a turn represents an uninterrupted, continuous utterance by a single speaker.
+Two dialogue sets were subsampled and de-identified: (1) 150 randomly sampled dialogues to assess the overall strategy alignment, and (2) one dialogue per week from 7 randomly selected participants to assess ChatWise’s alignment robustness to different participants and time periods over the clinical trial.
+4.1.2 Offline Evaluation Metric
+We define Strategy Match Percentage (SMP) to quantify the alignment between the conversational strategies generated by a chatbot and those generated by a human professional caregiver from the I-CONECT clinic data.
+Let $\mathbf{a}_{t}$ denote the set of caregiver-provided strategies annotated from a real conversation $\tau$ at turn (timestep) $t$, and $\mathbf{a}^{\prime}_{t}$ the strategies from a simulated moderator (ChatWise or utterance generator), conditioned on the same real context $s_{t}$.
+Let $\mathbb{I}(a^{\prime}_{t}\in\mathbf{a}_{t})$ be an indicator function that returns 1 if a chatbot-generated strategy $a^{\prime}_{t}$ matches one of the human strategies $\mathbf{a_{t}}$, and 0 otherwise.
+The SMP is computed as following:
+$\text{SMP}_{t}(\mathbf{a}_{t},\mathbf{a}^{\prime}_{t})=\frac{1}{|\mathbf{a}_{t}|}\sum_{a^{\prime}_{t}\in\mathbf{a}^{\prime}_{t}}\mathbb{I}(a^{\prime}_{t}\in\mathbf{a}_{t}).$
+Figure 3: Comparison of moderator strategies: Given real dialogue contexts, responses from a baseline (generated), ChatWise (generated), and the human moderator (real) are compared w.r.t.
+underlying strategies.
+A real sample is shown here while omitting participant data.
+4.1.3 Strategy Label Annotation
+Given a real dialogue denoting as $\tau=\{\mu_{0},u_{0},x_{1},u_{1},x_{2},u_{2},\cdots\}$ with caregiver (moderator) utterance $u_{t}$ and participant utterance $x_{t}$ at turn $t$, we leveraged the macro-level strategy definition in Sec 3.1 to prompt an LLM as annotator (GPT-4o) and detect strategies $\mathbf{a}_{t}$ behind $u_{t}$ as the golden label.
+Denoting $s_{t}=\{\mu_{0},u_{0},x_{1},\cdots,x_{t}\}$ as the real conversation context, we feed $s_{t}$ to a baseline chatbot to get utterance action $\tilde{u}_{t}$.
+The baseline is identical to the utterance generator $\pi_{u}$ of ChatWise yet does not receive strategy guidance from $\pi_{s}$.
+Similarly, we get strategy labels $\tilde{\mathbf{a}}_{t}$ behind $\tilde{u}_{t}$ using the above annotator.
+We then input $s_{t}$ to ChatWise’s strategy provider $\pi_{s}$ to get proposed strategy $\mathbf{a}^{\prime}_{t}\sim\pi_{s}(s_{t})$.
+We report the strategy alignment of $\text{SMP}_{t}(\mathbf{a}_{t},\mathbf{a}^{\prime}_{t})$ (w/ ChatWise) and $\text{SMP}_{t}(\mathbf{a}_{t},\tilde{\mathbf{a}}_{t})$ (w/o ChatWise) over different turns $t$ and average the result across dialogues.
+Figure 3 illustrates the process of collecting each moderator’s response based on real conversation history.
+4.1.4 Overall Strategy Alignment
+As shown in Figure 5, which averages results over 150 dialogues, ChatWise consistently achieves an SMP close to 1.0, indicating a strong alignment with the strategies employed by real caregivers throughout the conversation.
+In contrast, the baseline chatbot deviates from human behavior when lacking action guidance.
+Both chatbots used GPT-4o as the utterance generator $\pi_{u}$, and o3-mini was used for $\pi_{s}$ in ChatWise.
+We do not employ micro-level utterance similarity, as the I-CONECT interventions involve spontaneous daily conversations, rather than fixed question answering, unlike domains in programming or math.
+Whereas, on a macro level, ChatWise mostly mirrors the strategic choices of human professionals who are well trained to follow clinical protocols.
+To illustrate how our method behaves differently from the baseline, we collect the real dialogue context and different moderator responses where $\mathbf{a_{t}}=\mathbf{a}^{\prime}_{t}\neq\tilde{\mathbf{a}_{t}}.$ and show the discrepancy heatmap.
+Raw dialogue content is omitted to comply with the data usage agreement.
+Results in Figure 4 demonstrate that while “open question” is a commonly dominant strategy (forward DA) for both human and chatbot moderators, the primary strategy difference lies in how they lead into the question (e.g.
+“acknowledge”, “restatement”), which may in turn influence the specific question ultimately asked.
+Figure 4: Strategy distribution for dialogue samples where $\mathbf{a_{t}}=\mathbf{a}^{\prime}_{t}\neq\tilde{\mathbf{a}_{t}}.$ The vertical axis indicates ground-truth strategies $\mathbf{a_{t}}$ extracted from human professional utterances, and the horizontal axis denotes strategies $\tilde{\mathbf{a}}_{t}$ detected from the baseline chatbot.
+Figure 5: SMP between human caregiver and chatbots on real conversation data, where w/o ChatWise denotes the baseline without strategy guidance.
+Our method strongly aligns with human caregiver behavior.
+Robustness to Participant Heterogeneity:
+Figure 6 presents the SMP of ChatWise across dialogue turns among participants.
+ChatWise consistently outperforms the baseline and maintains notably higher SMP.
+These results demonstrate the robust alignment of ChatWise  with human caregiver strategies across diverse participants and their interaction contexts, which implies its potential for personalized dialogue systems.
+Figure 6: SMP broken down by participants over 40 conversation turns, where each color indicates a specific person.
+Solid lines represent ChatWise, and dashed lines denote the baseline.
+ChatWise consistently aligns to human caregivers given different participant contexts.
+Alignment Consistency Over Timeline:
+Figure 7 presents the SMP over time, where dialogues were sampled from different weeks over the clinical intervention.
+ChatWise maintains consistently high SMP (near 1.0) across all timesteps, where the baseline shows lower and unstable performance.
+This indicates ChatWise’s temporal robustness and its potential to facilitate the development of long-term cognitively supportive dialogues.
+Figure 7: Strategy alignment of ChatWise is consistent as real conversations progress over weeks.
+4.2 Evaluation with Digital Twins
+4.2.1 Conversation Generation
+While evaluation on real clinic data shows strong alignment between ChatWise and professional caregivers, the conversations are fixed, thus lacking indications whether and how ChatWise may influence user cognitive or emotional status during conversations.
+To enable controlled and responsive development of ChatWise prior to future human user studies, we used 9 digital twins provided by  Hong et al.
+(2024) as simulated users with different personas, which are fine-tuned LLMs calibrated to mimic the linguistic behaviors of older adults, some of whom show MCI symptoms.
+All digital twins are trained using real, de-identified dialogue data from the I-CONECT clinical trial.
+We collected 20 trajectories of conversations between ChatWise (moderator) and each digital twin as a user, with each conversation containing 20 turns.
+The same settings were applied to the baseline chatbot for fair comparisons.
+4.2.2 Metrics for Interactive Engagement
+User Verbosity: Given a conversation sequence $\tau$ that consists of user utterance $x$ and moderator utterance $u$, we primarily measure user engagement through the user’s talkativeness compared to the moderator, i.e.
+the user’s verbosity ($v$), defined as $v=\frac{\sum_{x\in\tau}|x|}{\sum_{u\in\tau}|u|},$ where $|x|$ ($|u|$) denotes the number of tokens in an utterance $x$ ($u$).
+This metric follows a clinical study (Yu et al., 2021) that suggests reducing the moderator talkativeness while encouraging participant (user) expression.
+Cognitive Win Rate:
+See et al.
+(2019) defined different aspects for evaluating conversation quality, from which we select 3 that are focused on assessing user cognitive status: Listening, Fluency, and Making Sense.
+All three metrics evaluate the user behavior rather than the overall dialogue quality considering both user and moderator behaviors.
+Instead of providing a numerical score for each cognitive metric that might be unstable, we adopt the Win Rate (WR) definition from prior work  (Rafailov et al., 2023) to compare pairs of dialogues generated with and without ChatWise.
+Dialogue pairs are randomly matched so long as they are collected from the same digital twin (simulated user) for evaluation.
+WR is defined as the proportion of pairs in which the dialogue generated with ChatWise is preferred over the baseline using an LLM-as-judge, based on a given cognitive metric.
+To mitigate the LLM’s position bias issues observed in prior work Shi et al.
+(2024); Yu et al.
+(2024), we excluded sample pairs with inconsistent preference labels when the order of the two samples was reversed and calculated WRs using only the remaining consistent pairs.
+Strategy Provider Verbosity $\uparrow$ Listening $\uparrow$ Fluency $\uparrow$ Making Sense $\uparrow$
+baseline 0.7398 0.5249 0.4986 0.5024
+GPT-4o 0.8635 0.4962 0.4748 0.4786
+o3-mini 0.8643 0.4884 0.5368 0.5180
+Llama 3.1-405B 0.8083 0.4407 0.4926 0.4963
+Table 1: Multifaceted evaluations on cognitive engagement with data sampled by interacting with digital twins over 20 turns, using different LLMs serving as strategy providers, where Listening, Fluency, and Making Sense metrics report win rates, and baseline denotes a conversation generator without receiving strategy guidance.
+All evaluations focus on user reactions.
+4.2.3 Models and Baseline
+We evaluated ChatWise with 3 different LLM backbones as strategy providers: GPT-4o, o3-mini, and Llama3.1-405B.
+For each setting, a baseline adopts the same utterance generator as ChatWise while without receiving strategy guidance.
+The WR of baseline is reported as the complement of ChatWise’s average WR: $\text{WR}_{\text{baseline}}=1-\text{average}(\text{WR}_{GPT-4o}+\text{WR}_{o3-mini}+\text{WR}_{Llama3.1-405B}),$ where $\text{WR}_{X}$ indicates the WR of ChatWise, given model $X$ as the strategy provider.
+4.2.4 Performance Overview
+Effectiveness of Strategy-Guided Generation:
+As shown in Table 1, our design mostly demonstrated better user engagement compared to baselines across tested LLMs as the strategy provider.
+Particularly, the best-performing model in our setting, o3-mini, was optimized for STEM reasoning tasks, which indicates that strong reasoning ability can also benefit inferring dialogue strategies.
+(a) User Verbosity.
+(b) WR: Listening.
+(c) WR: User Fluency.
+(d) WR: Making sense.
+Figure 8: Multifaceted Engagement Evaluation across conversation turns, excluding the first two warm-up turns. ChatWise’s performance gain over the baseline exists over multi-turn dialogues.
+Performance with Dialogue Progression:
+We analyzed ChatWise’s performance over increasing dialogue length by truncating conversations at turn $t$ and computing metrics accordingly, as shown in Figure 8.
+Results revealed a clear upward trend in verbosity, which indicates increased talkativeness of the user rather than the ChatWise, as opposed to the baseline.
+Meanwhile, the WR gains of ChatWise either remain stable or show a slight decline as conversations progress, which we infer are subject to the length bias of LLMs Dubois et al.
+(2024) as the overall number of tokens of ChatWise-generated dialogues tends to decrease in later turns.
 4.3 Ablation Study
-We evaluated ChatWise without user emotion in the strategy provider’s output (Table 2). Its consistent performance confirms that strategies are the primary driver of engaging conversations.
-Strategy Provider GPT-4o GPT-4o w/o emotion w/o ChatWise
+We evaluated ChatWise with and without user emotion in the strategy provider’s output.
+As shown in Table 2, where both strategy providers used GPT-4o as the backbone, removing emotion information led to only a slight drop of performance, while removing ChatWise entirely resulted in a much lower verbosity score, which highlights that strategy guidance itself is the key driver of user engagement.
+Method ChatWise ChatWise w/o emotion w/o ChatWise
 Verbosity 0.8635 0.8463 0.7398
-Table 2: User verbosity under different strategy providers. GPT-4o w/o emotion keeps the strategy provider with user emotion information removed.
+Table 2: Ablative study on the performance of user verbosity when removing different contextualized information. w/o emotion denotes the strategy provider without user emotion information as input.
 4.4 ChatWise’s Robustness to User Persona
-Figure 4 shows the significant gain of ChatWise interacting with all simulated users, indicating its robustness to support varying senior characteristics.
-Figure 4: Users’ dialogue verbosity (log-normalized, Appendix A.5) across 9 digital twins using different LLMs as strategy providers, compared with baseline (w/o ChatWise).
-4.7 Dialogue Analysis
-We identified 4 digital twins for which ChatWise was most effective and collected 40 additional dialogues from each to conduct a deeper analysis, with the main findings summarized below.
-Transient user emotions: We defined emotion transition triplets as $(\text{e}_{i},\text{s},\text{e}_{i+1})$, where $\text{e}_{i}$ ($\text{e}_{i+1}$) represents the user’s emotion at turn $i$ ($i+1$), and $s$ is the strategy provided by ChatWise at turn $i$.
+As shown in Figure 9, the gain of ChatWise interacting with 9 simulated users is consistently significant, in which the metrics were log-normalized.
+This indicates the robustness of ChatWise  given varying senior characteristics.
+Figure 9: User verbosity log-normalized across 9 digital twins.
+Given varying strategy provider backbones, our design consistently outperforms the baseline (w/o ChatWise) with non-strategic conversations.
+4.5 Analysis of User Emotion Transitions
+To analyze if ChatWise can support positive emotional shifts during conversations, we selected four digital twins and collected 40 additional dialogues from each for deeper analysis, with the main findings summarized below.
+Transient user emotions:
+We define an emotion transition triplet as $(\text{e}_{t},u_{t},\text{e}_{t+1})$, where $\text{e}_{t}$ ($\text{e}_{t+1}$) represents the user’s emotion at turn $t$ ($t+1$), and $u_{t}$ is the moderator utterance in between.
 We computed the average occurrence of emotion triplets and showed the top 15 most frequent in Figure 10.
-Most triplets reflect unchanged user emotions, indicating the challenge of detecting or influencing emotional shifts within a short turn.
-Long-term user emotions: We analyzed user emotion changes from the beginning to the end of dialogues and calculated averaged occurrence across digital twins (Figure 6).
+Most triplets reflect unchanged user emotions, which indicates difficulty in either influencing or detecting emotional shifts within a short turn.
+Figure 10: Average occurrence of each emotion transition triplet across the samples of each digital twin.
+User Emotions over Multi-Turn Conversations:
+We then analyzed user emotion changes from the beginning to the end of dialogues and calculated averaged occurrence across digital twins (Figure 11(b)).
 Over 48% users experienced emotional shifts post-dialogue, with significantly more positive changes after engaging with ChatWise.
-This implies that while transient emotions are difficult to track, strategic conversational support can improve user emotions over time.
-Predominant strategies: We identified the top 10 most frequent strategies across digital twins and found that predominant interaction strategies remained consistent across user types (See Appendix A.8).
-Particularly, Open Question, Statement-non-opinion, and Acknowledgment strategies dominate ChatWise driven conversations, suggesting their potential efficacy in fostering engaging dialogues.
-Figure 5: Average occurrence of each emotion transition triplet across the samples of each digital twin.
-Figure 6: Occurrence of user emotional changes. For instance, “neutral-joy” indicates user emotion transitioning from neutral (conversation begins) to joy (ends).
+This implies that while transient emotions are difficult to track, strategic conversational support may improve user emotions over multi-turn conversations.
+(a) Real user emotion transitions summarized from 150 dialogues in the I-CONECT study.
+(b) Digital twin emotion transitions summarized over 80 dialgoues.
+Figure 11: User emotion shifts in conversations between digital-twin and ChatWise (Figure 11(b)) align well with real-user emotion shifts in the clinical study (Figure 11(a)).
+More neutural-to-joy transitions were observed given ChatWise compared to the baseline.
+Predominant Strategies:
+We identified the top 10 most frequent strategies across digital twins (Figure 13) and found that predominant interaction strategies remained consistent across user types (See Appendix A.8).
+Particularly, Open Question, Statement-non-opinion, and Acknowledgment strategies dominate ChatWise driven conversations.
 5 Conclusion and Future Work
-We introduce ChatWise, a tuning-free yet principled framework for developing AI chatbots that foster long-term user engagement.
-Our dual-level generation design integrates clinical insights into LLM reasoning and is compatible with various LLMs.
-Extensive evaluations based on de-identified clinical data showed that our method robustly enhances the engagement and cognitive status of seniors.
-Moving forward, we aim to 1) enhance multi-turn interaction via RL and 2) validate its effectiveness through real user studies.
+We present a dual-level framework for AI chatbots that supports multi-turn conversations for older adults by integrating clinical insights into LLM reasoning.
+Extensive evaluations based on both real clinical data and generated conversations showed that our method aligns well with the behavior of professional human caregivers while robustly enhancing the cognitive status of older adults through simulation studies.
+Our future work includes training-based methods for further optimizing multi-turn interaction experience and investigating the acceptability and feasibility of our design through real user studies.
+6 Ethical Considerations
+This study involved simulated dialogues using digital twins trained on de-identified conversation data from a clinical trial involving older adults.
+Offline analysis of this dataset was conducted to evaluate the similarity between chatbot and caregiver strategies.
+This project did not use any personally identifiable information, and all data processing followed institutional privacy and research ethics guidelines.
+Potential risks are controllable, which include privacy concerns of emotion detection and misinterpretation of conversation due to AI-based evaluation.
+To mitigate this, the emotional detection module in the system is made optional and can be omitted as configured.
+Win rate calculations are based on a sanity check where we omitted pair samples that receive inconsistent preference labels by LLM-as-the-judge after swapping their positions.
+We have manually cross-validated a sufficient subset of AI-generated analyses to reduce bias.
+A real-user study is planned for future validation and will undergo full Institutional Review Board (IRB) approval.
 Limitations
-All digital twins were provided as fine-tuned GPT-3.5 APIs, and its high economic cost, at the time when this project was going, constrained the volume of dialogues we generated.
-To mitigate this, future work will explore training digital twins using LLaMA 3.1 to reduce sampling costs while maintaining conversational quality.
-Additionally, our study relies on simulated dialogues rather than real user interactions.
-A real user study is needed to validate the system’s effectiveness in dynamic settings, which we plan to incorporate into our future research.
+All digital twins were provided as fine-tuned GPT-3.5 APIs, and their high economic cost, at the time when this project was going on, constrained the volume of dialogues we generated.
+To mitigate this, future work will explore accessing digital twins using LLaMA 3.1 to reduce sampling costs while maintaining conversational quality.
+LLM evaluations have been discussed to show position bias.
+Regarding win rate calculation, after we took mitigation methods and filtered out pairs with inconsistent preference labels, the remaining samples contributed to 81% of the data before filtering.
+Future mitigation methods will include learning a reward function for more fine-grained and robust evaluation.
+Furthermore, our study relies either on offline static data or simulated dialogues rather than dynamic user interactions.
+A real user study is needed to validate the system’s acceptability and feasibility in real-world settings, which we plan to incorporate into our future research.
 Appendix A Appendix
 A.1 Prompt design
-The prompt used in the experiments and structured output design are available at https://anonymous.4open.science/r/ChatWise-D5E0, including:
+The prompt used in the experiments and structured output design are available at https://anonymous.4open.science/r/ChatWise-8F53, including:
   • Strategy provider system prompt.
@@
   • System prompt for GPT-4o to extract the strategy given by Llama3.1.
-A.2 Data Generation Configuration
-We used GPT-4o as the utterance generator in ChatWise and tested different LLM backbones as strategy providers.
+A.2 Offline Data Preparation
+We extracted dialogue content from the I-CONECT video conversations and processed it into OpenAI-compatible dialogue history format.
+Dialogues with fewer than 40 turns were excluded.
+We then subsampled two subsets for evaluation purposes.
+To assess the overall strategy alignment of ChatWise, we randomly sampled 150 dialogues.
+To evaluate the robustness of ChatWise across different participants and time periods, we randomly selected 7 participants who completed the full study and sampled one dialogue from each of their sessions during weeks 1, 9, 17, 25, 33, and 41.
+All user identifiers and personally identifiable information have been anonymized.
+The OpenAI-compatible dialogue history format example:[{"role": "system", "content": system_prompt}, {"role": "user", "content": user_content1}, {"role": "assistant", "content":assistant_content1}, {"role": "user", "content": user_content2}, {"role": "assistant", "content": assistant_content2}, …].
+A.3 Data Generation Configuration
+By default, we used o3-mini as strategy provider.
+GPT-4o serves as an utterance generator in all settings.
+W/o ChatWise denotes the baseline, which is using GPT-4o as the utterance generator only, without a strategy provider.
+We tested different LLM backbones as strategy providers in ChatWise.
 Considering the Llama3.1-405B does not support structured output, we applied GPT-4o as the strategy extractor to structure the strategy output of Llama3.1-405B.
+We employed GPT-4o as the judge to select the preferred one when computing the Win Rate.
+The prompt for selecting the preferred response is listed in Figure 12.
 The following are the configurations of each model:
@@
 temperature=1
-A.3 Win rate for w/o ChatWise
+GPT-4o as judge:
+n=1
+max_tokens=1024
+top_p=1
+temperature=1
+Figure 12: Prompt for GPT-4o as a Judge.
+A.4 Win rate for w/o ChatWise
 The Win rate for w/o ChatWise is defined as:
@@
 ,where $\text{WR}_{X}$ is ChatWise’s Win rate against the baseline with $X$ as strategy provider.
-A.4 Dialogue Acts
+A.5 Dialogue Acts
 The map of strategy to its corresponding abbreviated tag is listed in Table3.
+There are two kinds of strategies: backward-looking and forward-looking.
+Backward-looking strategies reflect how the current utterance relates to the previous discourse.
+Forward-looking strategies reflect the current utterance constrains the future beliefs and actions of the participants and affects the discourse.
+Table 4 and Table 5 provide definitions and examples of each.
+A.6 Log-normalization
+The following is the log-normalization function, where y is the normalized result, x is the input variable.
+$y=\ln{\left(4x+1\right)}$
+A.7 Primary Strategies
+We calculated the average occurrence of each strategy across each digital twin and listed their top 10 most frequently occurring strategies, as shown in Figure 13.
+Figure 13: Strategy occurrence across digital twins.
 Strategy Tag
@@
 Information I
-Table 3: Dialogue Act to its corresponding tag.
-There are two kinds of strategies: backward-looking and forward-looking.
-Backward-looking strategies reflect how the current utterance relates to the previous discourse.
-Forward-looking strategies reflect the current utterance constrains the future beliefs and actions of the participants and affects the discourse.
-Table 4 and Table 5 provide definitions and examples of each.
+Table 3: Strategy to its corresponding tag.
+The strategies are drawn from the DAs in DAMSL  (Allen and Core, 1997) that are used by telehealth clinical trials  (Yuan et al., 2023), integrated strategies from prior emotional support dataset  (Liu et al., 2021) .
 Strategy Definiton Example
@@
 Table 5: Forward-looking strategies, definition, and example.
-A.5 Log-normalization
-The following is the log-normalization function, where y is the normalized result, x is the input variable.
-$y=\ln{\left(4x+1\right)}$
-A.6 Additional results for Accumulated Multi-Turn Analysis
-ChatWise’s accumulated performance on Fluency, and Making Sense is shown in Figure 7.
-(a) Fluency Across Turns
-(b) Making sense Across Turns
-Figure 7: ChatWise performance with o3-mini as strategy provider across turns. W/o ChatWise is the baseline that ChatWise is not applied. GPT-4o servers as moderator in all settings.
-A.7 Primary Strategies
-We calculated the average occurrence of each strategy across each digital twin and listed their top 10 most frequently occurring strategies, as shown in Figure 8.
-Figure 8: Strategy occurrence across digital twins.
 A.8 Personalized Dialogue Analysis